OpenStudy (dls):

In a chemical reaction,the rate constant of backward reaction is 7.5 x 10^-4 and the eqm. constant K=15..What is the rate of forward reaction?

OpenStudy (dls):

\[\LARGE 1.5=\frac{7.5 \times 10^{-4}}{x}\]

OpenStudy (dls):

Is that how it will be done ?

OpenStudy (shubhamsrg):

LHS = 1/1.5

OpenStudy (dls):

??

OpenStudy (shubhamsrg):

1.5 = x/(7.5)

OpenStudy (shubhamsrg):

*10^something

OpenStudy (dls):

what is the formula?

OpenStudy (dls):

\[\LARGE K_c = \frac{Backward}{Forward}\]?

OpenStudy (shubhamsrg):

K = product conc./react conc.

OpenStudy (dls):

same thing :o

OpenStudy (dls):

i meant that only :P

OpenStudy (shubhamsrg):

product conc directly proportional to forward reac same for reactant conc

OpenStudy (dls):

?

OpenStudy (dls):

rate constant and eqm constant=same?

OpenStudy (shubhamsrg):

eqbm const = forward rate/backward rate

OpenStudy (dls):

:/

OpenStudy (dls):

\[\LARGE A+B \rightarrow C+D\] to [C][D]/[A][B] ? :/

OpenStudy (shubhamsrg):

ha

OpenStudy (dls):

to back/for?

OpenStudy (shubhamsrg):

hein ?

OpenStudy (dls):

ok nvm thanks

OpenStudy (shubhamsrg):

thanks? :O

OpenStudy (shubhamsrg):

pehli baar! :')

OpenStudy (dls):

not fully satisfied,fact samaj ke maannlia isiliye that sarcastic thx

OpenStudy (shubhamsrg):

samjh kyu nahi aya? o.O

OpenStudy (dls):

ratlia bus

OpenStudy (shubhamsrg):

C+D ban na , matlab forward reacn

OpenStudy (dls):

:O

OpenStudy (dls):

le yaar :P

OpenStudy (shubhamsrg):

:'D

OpenStudy (dls):

:|